Write the net ionic equation for the precipitation reaction that occurs when aqueous solutions of aluminum acetate and ammonium carbonate are combined.

Use the pull-down boxes to specify states such as (aq) or (s).
Answer #1

The complete molecular equation is:

2 Al (OCOCH3)3 (aq) + 3 (NH4)2CO3 (aq) = Al2 (CO3)3 (s) + 6 NH4OCOCH3

Net ionic equation :

2 Al+3 (aq) + 3 CO3-2 (aq) = Al2 (CO3)3 (s)